OpenStudy (anonymous):

What is the approximate volume of a cone with a height of 5 meters and base radius of 8 meters? if you can, please use 3.14 for pi two hundred fifty-one point two cubic meters three hundred thirty-four point nine cubic meters six hundred twenty-eight cubic meters six hundred fifty-three point one two cubic meters

OpenStudy (anonymous):

Ok, which equation do we need to use to find the volume of a cone?

OpenStudy (anonymous):

SA =r2 +rt I think?

OpenStudy (anonymous):

when pi (about 3.14) r: radius of circle t: slant height of cone

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

What does SA stand for...surface area? I think were trying to solve for volume here....correct me if I'm wrong.

OpenStudy (anonymous):

0-0 oh yeah, sorry its 1/3 * 3.14 * r^2 * h with r being the radius, and h being height right?

OpenStudy (anonymous):

Yup, so now we just need to put in what we know. We know the radius is 8, and we know the height is 5, so we can solve this. Try putting in the numbers ans see if you can figure it out yourself.

OpenStudy (anonymous):

is it 628?

OpenStudy (anonymous):

Hmm..that's not what I got... Here's what I did: V = 1/3*3.14*(8)^2*5 V = 1/3*3.14*(64)*5 V= 1/3*3.14*320 V = 1/3*1004.8 V = 334.933333 rounds down to V = 334.9 cubic meters. Double check your work maybe? :o

OpenStudy (anonymous):

oh ok, sorry, I added 3.14 by 3.14 XD

OpenStudy (anonymous):

Not sure why you did that. Lol. I believe the answer is B then. :)